OpenStudy (anonymous):

List the three methods for solving equations and give a brief explanation of each method.

hero (hero):

Three methods for solving SYSTEMS of equations?

OpenStudy (anonymous):

Yes

hero (hero):

Hang on

OpenStudy (anonymous):

Ok

hero (hero):

1. By Elimination: If you have a system of the form: 3x `+ 2y` = 4 4x `- 2y ` = 10 you notice that 2y and -2y are opposites, then you combine the equations together to eliminate the y variable. When you do that, you get: 7x = 14 Solving for x to get x = 2, which means y = -1 2. By substitution If you have a system of the form: x = `2y + 4` 2x - y = 8 Then you notice that one of the variables is isolated. In this case, x is isolated in the first equation, therefore you substitute the expression for x into the second equation to get: 2( `2y + 4`) - y = 8 Then solve for y: 4y + 8 - y = 8 4y - y + 8 = 8 3y = 8 - 8 3y = 0 y = 0/3 y = 0 So y = 0. Substituting y back in to the original equation yields x = 4 By graphing: If you have a system of the form y = 2x + 4 y = 7x - 3 You can solve the system by graphing since both equations are in the form y = mx + b to find the solution point (x,y).

OpenStudy (anonymous):

Thank You!